OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. A.y = x + 2 B.y = -x - 2 C. y = -x + 4 D. y = x - 4

OpenStudy (mathstudent55):

Parallel lines have the same slope. When you write the eq of a line as y = mx + b, m is the slope. This eliminates two answers.

OpenStudy (mathstudent55):

Since the given line has a slope od -1, the parallel line has to be y = -x + b. To find b, insert the values of x and y of the given point into y = -x + b and solve for b.

OpenStudy (anonymous):

i dont under stand

OpenStudy (mathstudent55):

Slope is a measure of the inclination of a line. It tells you how much the line tilts one way or the other. When you graph a line, a line can be horizontal, vertical, or diagonal. When it is diagonal, it either goes up to the right or up to the left.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The slope tells you exactly how th eline tilts. If two lines are parallel, they have the same slope.

OpenStudy (mathstudent55):